### v5

# On Picard’s Theorem Via Nevanlinna Theory

Abstract.
We consider the Picard’s problem via Nevanlinna theory for complete Kähler manifolds with non-negative Ricci curvature. When the manifold is non-parabolic, which means that it carries a positive global Green function, we give a positive answer to the Picard’s problem. That is, we confirm that every meromorphic function on such a (non-compact) manifold reduces to a constant if it omits three distinct values.
Key words and phrases:
Nevanlinna theory; defect relation; Picard type theorem2010 Mathematics Subject Classification:
32H30; 32H25Contents
1. Introduction
1.1. Motivations and Techniques
It is well-known that the famous Picard’s little theorem asserts that every meromorphic function on the complex Euclidean space reduces to a constant, if it omits three distinct values. All the time, many authors have been trying to generalize this theorem to a more general non-compact complex manifold. Due to S. T. Yau [44], each holomorphic function on a non-compact complete Kähler manifold (or Hermitian manifolds in general) with non-negative Ricci curvature has Liouville property. So, a natural problem is that does Picard’s little theorem still hold on such a manifold? We call it the Picard’s problem. To our knowledge, it is still an unsolved problem. In what follows, we would like to give a precise statement.
Let be a non-compact complete Kähler manifold with non-negative Ricci curvature, see examples for such a manifold in Sha-Yang [34] and Tian-Yau [39, 40]. The well-known Picard’s problem asks that
Picard’s Problem. Is every meromorphic function on necessarily a constant if it omits distinct values
Remark. Let be the universal covering, where is the unit disc in It is not very hard to confirm that the Picard’s problem can be given a positive answer using Yau’s theorem (see Theorem C below), if any holomorphic mapping lifts to a holomorphic mapping via i.e., Using the criterion for the existence of the liftings, can be lifted to if and only if for some such that Since is simply-connected, then is trivial. This implies that such a lifting exists if and only if is trivial. However, is often not a trivial group, so one cannot lift to in general.
The first result in this direction could be traced back to 1970, S. Kobayashi [20] proved some Picard-type theorem for a holomorphic mappings between complex manifolds from the viewpoint of complex hyperbolicity. As a special consequence, he showed that
Theorem A (Picard’s Theorem, Kobayashi).
Every meromorphic function on on which, a complex Lie group acts transitively, must be a constant if it omits distinct values.
Kobayashi’s result is limited to certain complex manifolds which are acted on transitively by a complex Lie group. In 1975, Goldberg-Ishihara-Petridis [17] presented some Picard-type theorems on harmonic mappings of bounded dilatation on a locally flat Riemannian manifold, refer to N. Petridis [27] also. In particular, they showed that
Theorem B (Picard’s Theorem, Goldberg-Ishihara-Petridis).
Assume that is locally flat. Every holomorphic mapping of bounded dilatation must be a constant mapping.
It does not seem to be a quite satisfactory answer to the Picard’s problem, for some restrictions such as local flatness and bounded dilatation. The first result without any restrictions for is due to S. T. Yau [44], who considered the Liouville’s problem based on the gradient estimation method at the same time, which is viewed as a weaker version of the Picard’s problem. He showed that
Theorem C (Liouville’s Theorem, Yau).
Every holomorphic function on must be a constant if it is bounded.
By S. T. Yau [44], Liouville’s theorem works for all the harmonic functions on (see Cheng-Yau [9] also). The gradient estimation method shows great power in many problems from geometric analysis (see, e.g., [35]), but it seems to be unable to give a solution to the Picard’s problem. Until 2010, A. Atsuji [2] developed a Nevanlinna-type theory based on heat diffusions, and he gave a positive answer to the Picard’s problem for as a meromorphic function on is of a slow growth rate. In the following, we introduce his basic ideas. Let be the Kähler form and Laplace-Beltrami operator of associated to respectively. One can endow with Fubini-Study metric Let be a meromorphic function. Using the Kählerness of the Hilbert-Schmidt norm of differential (with respect to metrics ) can be written as the form
He showed that
Theorem D (Picard’s Theorem, Atsuji).
Let be a meromorphic function. Assume that satisfies the growth condition
where is a geodesic ball centered at a fixed point with radius in Then, must be a constant if it omits distinct values.
Atsuji’s trick [2] (see [11] also) is employing the probabilistic approach (via Brownian motions), who introduced the so-called Nevanlinna-type functions and in which is the time of a Brownian motion on By using Itô’s formula (see, e.g., [18]) and estimates of curvatures, he established a heat diffusion version of Second Main Theorem, which leads to the above Theorem D. It is apparent that is stochastically complete (or is conservative) since has non-negative Ricci curvature by Grigor’yan’s criterion [15]. However, in order to make and meaningful in spirit of Nevanlinna’s settings, the following conditions are necessary:
for ;
as ;
if omits
To do so, has to satisfy certain growth assumptions. That is why a growth condition is needed in Atsuji’s theorem. Unfortunately, the growth condition is so strong that few meromorphic functions can satisfy it.
In the present paper, we shall consider the Picard’s problem by developing Nevanlinna theory (see, e.g., [24, 26, 29]) to complete Kähler manifolds with non-negative Ricci curvature. As we know, Nevanlinna theory studies value distribution of meromorphic mappings between complex spaces, which has been developed for a long time since R. Nevanlinna founded two fundamental theorems for meromorphic functions on the complex plane in 1925. During this period, rich results have been achieved, we may refer the reader to L. V. Ahlfors [4], H. Cartan [6], Carlson-Griffths-King [7, 16], J. Noguchi [23, 24], E. I. Nochka [25], M. Ru [28, 29], B. Shiffman [30], B. Shabat [31], F. Sakai [32, 33], W. Stoll [37, 38], P. Vojta [36], H. Wu [43] and etc., and refer also to [1, 2, 3, 11, 12, 13] and references therein. To our knowledge, almost all the results in Nevanlinna theory are built just on Kähler manifolds which admit a complete Kähler metric of non-positive sectional curvature, while we seem to know little if the domain manifolds are not non-positively curved, because it is hard to estimate the local Green functions for bounded domains under other curvature conditions.
Recently, the author [14] obtained the first result for the value distribution of meromorphic mappings on a complete Kähler manifold with non-negative Ricci curvature by introducing a new method: global Green function method. Based on this method, a positive answer to the Picard’s problem can be given provided that is of maximal volume growth, i.e.,
where denotes the Riemannian volume of geodesic ball centered at a fixed reference point with radius in
The original technique in [14] is to construct a family of relatively compact domains exhausting by using a minimal positive global Green function for Based on the asymptotic estimate of obtained by Colding-Minicozzi [8], we can give an optimal estimate for the local Green function for in terms of integral forms. As a result, the author established the Second Main Theorem of meromorphic mappings on Let us introduce the main contribution and key technique in [14]. Let be the minimal positive global Green function of for Colding-Minicozzi [8] (see [21] also) described the asymptotic behavior of : there exists a constant such that
where denotes the Riemannian distance function of from By means of this asymptotic estimate, the author defined a family of relatively compact domains:
Thus, the Green function of for with a pole at satisfying Dirichlet boundary condition can be written as
According to Lemma 3.5 in [14], for any there exists such that
holds for all satisfying Based on this estimate, the Calculus Lemma can be derived. When is conformally equivalent to since has non-negative Ricci curvature and maximal volume growth (see [21]). So, by employing the standard arguments, the author established the Second Main Theorem of meromorphic mappings from into a complex projective manifold (see [14], Theorem 3.10), which gives a Picard’s theorem:
Theorem E (Picard’s Theorem, Dong).
Assume that is of maximal volume growth. Every meromorphic function on must be a constant if it omits distinct values.
However, the technique in [14] depends on Colding-Minicozzi’s asymptotic estimate. Unfortunately, there are no asymptotic estimates for if is not of maximal volume growth.
In this paper, we will consider a weaker volume growth condition:
which is equivalent to that carries a positive global Green function. Refer to N. Varopoulos [41, 42] (or see Li-Tam-Wang [19, 21]), there exists uniquely a minimal positive global Green function of for denoted by Our original method is the construction of domain and the application of Li-Yau’s estimates on heat kernels and Green functions.
By Li-Yau’s estimate [22], there exist constants such that
Re-define by
In further, we have
which defines the Green function of for with a pole at satisfying Dirichlet boundary condition. There is a natural relationship between and (see Lemma 4.1): for we have
which plays an important role in the establishment of Nevanlinna theory in this paper.
In the study of Nevanlinna theory, Negative Curvature Method and Logarithmic Derivative Lemma Method are considered as two standard research approaches. In our investigations, Logarithmic Derivative Lemma Method is used. In doing so, it is necessary to establish the so-called Calculus Lemma, which turns out to be a workable road to the Logarithmic Derivative Lemma.
Let us introduce the main techniques: Gradient Estimation (see Theorem 4.2) and Calculus Lemma (see Theorem 5.2) as follows.
Gradient Estimation. For any there exists such that
holds for all
The above gradient estimate of gives an estimate of the harmonic measure on with respect to : for any there exists such that (see Corollary 4.3)
holds for all
Calculus Lemma. Let be a locally integrable function on Assume that is locally bounded at Then for any there exist a constant and a subset of finite Lebesgue measure such that
holds for all outside
Let be a nonconstant meromorphic function on The Calculus Lemma can lead to Logarithmic Derivative Lemma (see Theorem 5.6): for any there exists a subset of finite Lebesgue measure such that
holds for all outside
1.2. Main results
Let be a complex projective manifold with over which one can put a Hermitian positive line bundle with Chern form Fix a reduced divisor where denotes the complete linear system of Given a meromorphic mapping we have the Nevanlinna’s functions and see definition in Section 3.2 of this paper. Assume that is non-parabolic, by which we mean that carries a positive global Green function, i.e., it satisfies the volume growth condition
Let be the Ricci form of associated with the metric and be the canonical line bundle over The characteristic function of is defined by
The main result is the following Second Main Theorem.
Theorem 1.1 (=Theorem 6.1).
Let be a non-compact complete Kähler manifold with non-negative Ricci curvature. Let be a complex projective manifold of complex dimension not greater than that of Let be a reduced divisor of simple normal crossing type, where is a positive line bundle over Let be a differentiably non-degenerate meromorphic mapping. If is non-parabolic, then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
For a divisor the simple defect of with respect to is defined by
Again, put
Theorem 1.1 gives a defect relation:
Corollary 1.2 (=Corollary 6.4).
Assume the same conditions as in Theorem Then
Let denote the scalar curvature of Set
We obtain a Picard’s theorem:
Corollary 1.3 (=Corollary 6.11).
Every meromorphic function on must be a constant if it omits distinct values, where denotes the maximal integer not greater than
Corollary 1.3 shows that each meromorphic function satisfying on must be a constant, if it omits 2 distinct values.
2. Some Facts from Differential Geometry and Geometric Analysis
2.1. Volume Comparison Theorem
A space form is defined as a complete Riemannian manifold with constant sectional curvature. Given a simply-connected space form with constant sectional curvature of dimension Let denote the Riemannian volume of a geodesic ball with radius in Besides, let be a complete Riemannian manifold with Ricci curvature of dimension Fix a point Let denote the Riemannian volume of a geodesic ball centered at with radius
The classical volume comparison theorem by Bishop-Gromov (see [5]) can be stated as follows.
Theorem 2.1 (Volume Comparison Theorem).
If for a constant then the volume ratio is non-increasing in and it tends to as Hence, we have
holds for all
In particular, if then we obtain:
Corollary 2.2.
Assume that is non-compact. If then we have
holds for all where is the volume of a unit ball in
When Calabi-Yau (see [35]) obtained a lower bound of
Theorem 2.3.
Assume that is non-compact. If then has an infinite volume. More precisely, for any there exists a constant such that
holds for all
2.2. Estimates of Heat Kernels
Let be a complete Riemannian manifold of dimension with Laplace-Beltrami operator The heat kernel of is the minimal positive fundamental solution of the following heat equation
Fix any reference point Let be the Riemannian distance function of from
Theorem 2.4.
Assume that has non-negative Ricci curvature. Then for any there exist constants and such that
holds for all
Set
If the right-hand integral converges , then defines a minimal positive global Green function of for with a pole at i.e.,
where is the Dirac’s delta function with a pole at It is evident that such a Green function is also unique if (1) is satisfied.
Theorem 2.5.
Assume that has non-negative Ricci curvature. If exists, then there exist constants depending only on such that
holds for all
3. Nevanlinna’s Functions and First Main Theorem
Let be a non-compact complete Kähler manifold with non-negative Ricci curvature of complex dimension whose Kähler form is defined by
in a local holomorphic coordinate Fix a reference point Let stand for the geodesic ball centered at with radius in Denote by the Riemannian volume of
3.1. Construction of
Assume that is non-parabolic, which means that satisfies the volume growth condition
Hence, there exists a unique minimal positive global Green function of for i.e.,
where denotes the Laplace-Beltrami operator and is the Dirac function with a pole at Let be the Riemannian distance function of from According to Li-Yau [22], there exist constants such that
Define
By as it is evident to see that is relatively compact for any and the sequence exhausts for According to Sard’s theorem, the boundary of is a submanifold of for almost every Set
which is the positive Green function of for with a pole at satisfying Dirichelet boundary condition, i.e.,
Moreover, we denote by the harmonic measure on with respect to which is defined by
where is the inward normal derivative on is the Riemannian area element of
3.2. Nevanlinna’s Functions
In what follows, we will introduce Nevanlinna’s functions. Let be a meromorphic mapping, where is a complex projective manifold. Let be a Hermitian holomorphic line bundle over with the Chern form where
Fix a divisor where is the complete linear system of Let be the canonical section associated to i.e., is a holomorphic section of over with zero divisor The characteristic function, proximity function, counting function and simple counting function of are respectively defined by
where is the Riemannian volume element of
Equip with the Kähler metric Let be the Hilbert-Schmidt norm of differential Since is Kählerian, then we have
Locally, write where is a local holomorphic frame of and is a holomorphic function. By Poincaré-Lelong formula (see, e.g., [7]), it leads to
in the sense of currents. So, we obtain the alternative expressions of and as follows
and
3.3. First Main Theorem
To establish the First Main Theorem of we need Jensen-Dynkin formula (see, e.g., [12, 13]) as follows.
Lemma 3.1 (Jensen-Dynkin formula).
Let be a -class function on outside a polar set of singularities at most. Assume that Then
Assume that Apply Jensen-Dynkin formula to then we are led to
Therefore, we are led to that
Theorem 3.2 (First Main Theorem).
Assume that Then
4. Gradient Estimates of Green Functions
Let be a non-compact complete Kähler manifold of complex dimension with non-negative Ricci curvature. Assume that is non-parabolic.
Theorem 4.1.
We have
holds for all where is given by
Proof.
According to the definition of Green function for it is immediate that for
Since
then we obtain
Let denote the gradient operator on We obtain an estimate of upper bounds of as follows:
Theorem 4.2.
For any there exists such that
holds for all where is given by
Proof.
It yields from (2) that
On the other hand, we have
Whence, it concludes that
where is the cut locus of and is the inward normal derivative on By
we see that for any there exists such that
holds for all Since has measure 0, with the aid of the continuity of on and on one then has the theorem proved. ∎
As a result, we obtain an estimate for upper bounds of harmonic measure on with respect to as follows:
Corollary 4.3.
For any there exists such that
holds for all where is given by
Proof.
5. Calculus Lemma and Logarithmic Derivative Lemma
Let be a non-compact complete Kähler manifold of complex dimension with non-negative Ricci curvature. Assume that is non-parabolic.
5.1. Calculus Lemma
We need the following Borel’s lemma (see, e.g., [24]):
Lemma 5.1 (Borel Lemma).
Let be a non-decreasing function on with Then for any there exists a subset of finite Lebesgue measure such that such that
holds for all outside
Proof.
The conclusion is true clearly for Next, we assume that Since is a non-decreasing function, then there exists a number such that The non-decreasing property of implies that the limit
exists, here is allowed. If then Set
Note that exists for almost every Then, we have
This completes the proof. ∎
Now, we give the Calculus Lemma:
Theorem 5.2 (Calculus Lemma).
Let be a locally integrable function on Assume that is locally bounded at Then for any there exist a constant and a subset of finite Lebesgue measure such that
holds for all outside
Proof.
Invoking Theorem 4.1, we obtain
Set
A simple computation leads to
In further, we have
Apply Borel’s lemma to the left hand side of the above equality twice: one is to and the other is to we see that for any there exists a subset of finite Lebesgue measure such that
holds for all outside On the other hand, Corollary 4.3 implies that there exists such that
holds for all where is given by (2). Combining the above, we have
holds for all outside Therefore, we have the theorem proved by setting
As a result, we conclude that
Corollary 5.3.
Let be a locally integrable function on Assume that is locally bounded at Assume that is of maximal volume growth. Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
Proof.
5.2. Logarithmic Derivative Lemma
Let be a meromorphic function on The norm of the gradient is defined by
in a local holomorphic coordinate where is the inverse of Again, define the Nevanlinna’s characteristic function of by
where
It is not difficult to show that
On one puts a singular metric
Then, it leads to
Lemma 5.4.
We have
Proof.
Note that
Whence, it concludes from Fubini’s theorem that
Lemma 5.5.
Let be a meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
Proof.
Define
Next, we establish the following Logarithmic Derivative Lemma:
Theorem 5.6 (Logarithmic Derivative Lemma).
Let be a meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
Proof.
Note that
Using Lemma 5.5 and disturbing we have for any there exists a subset of finite Lebesgue measure such that
holds for all outside ∎
6. Second Main Theorem and Defect Relation
Let be a non-compact complete Kähler manifold of complex dimension with non-negative Ricci curvature. Assume that is non-parabolic.
Set
where is the Ricci form of
Let be a divisor, where are prime divisors. The reduced form of is defined as
6.1. Second Main Theorem
Let be a reduced divisor of simple normal crossing type. We write which is an irreducible decomposition of Equip every holomorphic line bundle with a Hermitian metric which induces a Hermitian metric on Since then we can assume that As a result, gives a volume form on Pick such that and Moreover, we define a singular volume form
on Let
be a differentiably non-degenerate meromorphic mapping. Set
It is clear that
In further, we have
in the sense of currents, where is the reduced form of Thus, it yields that
We establish a Second Main Theorem as follows.
Theorem 6.1 (Second Main Theorem).
Let be a non-compact complete Kähler manifold with non-negative Ricci curvature. Let be a complex projective manifold of complex dimension not greater than that of Let be a reduced divisor of simple normal crossing type, where is a positive line bundle over Let be a differentiably non-degenerate meromorphic mapping. If is non-parabolic, then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
Proof.
The argument of proof is standard (see [14]), but for the completeness of this paper, we still give a proof. Since has only simple normal crossings, then there exist a finite open covering of and finitely many rational functions on such that are holomorphic on satisfying that
In addition, we can require that for On write
where is a positive smooth function on Let be a partition of the unity subordinate to Set
Again, put On we have
Fix a one can take a local holomorphic coordinate near and a local holomorphic coordinate near such that
Put
Then, we have and at :
Define a non-negative function on by
Again, set for Then, at :
That is, at :
Combining the above, we are led to
on Note that is bounded on whence it yields from that
By Jensen-Dynkin formula
Combining (6) with (7) and using Theorem 5.6, we obtain
Using Theorem 5.2 and (5), for any there exists a subset of finite Lebesgue measure such that
holds for all outside Thus, we conclude that
for all outside Combining this with (6.1), we prove the theorem. ∎
Let be a compact Riemann surface of genus Let be the holomorphic line bundle over defined by a point Note that is independent of the choice of It yields from that
Corollary 6.2.
Let be distinct points in Let be a nonconstant meromorphic mapping. Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
Let be the hyperplane line bundle over and whose Chern form is given by It yields from that
Corollary 6.3.
Let be hyperplanes in general position in Let be a differentiably non-degenerate meromorphic mapping. Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
6.2. Defect Relation
For any divisor the defect and simple defect of with respect to are defined respectively by
Using the First Main Theorem, we see that
For two holomorphic line bundles over we define (see [7])
Corollary 6.4 (Defect Relation).
Assume the same conditions as in Theorem Then
Endowing with Kähler metric Let be the Hilbert-Schmidt norm of differential with respect to metrics Let denote the scalar curvature of Take a local holomorphic coordinate the Ricci curvature tensor of can be written as the form
where
Using the definition of scalar curvature, we have
In further, it yields that
Corollary 6.5.
Let be a meromorphic mapping. If there exists a small number such that satisfies the growth-curvature condition
for sufficiently large then must be differentiably degenerate.
Proof.
Assume on the contrary that is differentiably non-degenerate. Then, we have as The growth-curvature condition leads to
Thus, we obtain due to Corollary 6.4, but which contradicts with ∎
Treat a compact Riemann surface of genus we have:
Corollary 6.6.
Let be a nonconstant meromorphic mapping. Let be distinct points in Then
For a meromorphic mapping
we have
with respect to metrics
Corollary 6.7.
Let be a differentiably non-degenerate meromorphic mapping. Let be hyperplanes in general position in Then
A meromorphic function on can be seen as a meromorphic mapping Set
We obtain a Picard’s theorem:
Corollary 6.8 (Picard’s Theorem).
Every meromorphic function on must be a constant if it omits distinct values, where denotes the maximal integer not greater than
6.3. Picard’s Theorem
In the following, we investigate a more exact form of the Picard’s theorem given in Corollary 6.8. On we introduce the spherical distance
Let be a meromorphic function on which is regarded as a meromorphic mapping
In terms of the spherical distance, the proximity function of with respect to can be defined as
We use the Ahlfors-Shimizu’s characteristic function of :
Moreover, we define
It is not hard to show the First Main Theorem:
Theorem 6.9 (Second Main Theorem).
Let be a non-compact complete Kähler manifold with non-negative Ricci curvature. Let be distinct values in Let be a nonconstant meromorphic function on If is non-parabolic, then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
Proof.
By
in the sense of distributions, we deduce that
Combining this with Corollary 6.3, it gives immediately
We re-define the simple defect of with respect to by
Set
Corollary 6.10 (Defect Relation).
Assume the same conditions as in Theorem Then
Consequently, we derive a more exact Picard’s theorem:
Corollary 6.11 (Picard’s Theorem).
Every meromorphic function on must be a constant if it omits distinct values, where denotes the maximal integer not greater than