### v6

# On Picard’s Theorem Via Nevanlinna Theory

Abstract.
We study the classical Picard’s problem for complete Kähler manifolds with non-negative Ricci curvature. Based on the global Green function method, we generalize the Nevanlinna theory to complete Kähler manifolds with non-negative Ricci curvature. As a consequence, we give a positive answer to the Picard’s problem in the non-parabolic case, i.e., if the (non-compact) manifold is non-parabolic, then we show that every meromorphic function is a constant if it omits three distinct values.
Key words and phrases:
Nevanlinna theory; meromorphic mappings; Picard’s theorem; defect relation; Kähler manifolds2010 Mathematics Subject Classification:
32H30; 32H25; 32A22Contents
1. Introduction
1.1. Motivations and Techniques
The famous Picard’s theorem asserts that a meromorphic function on the complex Euclidean spaces will reduce to a constant, if it omits three distinct values. Historically, people have been committed to developing this theorem for a long time (see [4, 16, 18, 20, 21, 24, 31]). The study of Picard’s problem on manifolds may have originated from the celebrated work due to S. T. Yau [48] on Liouville’s theorem, which proves the Liouville property of harmonic functions on non-negatively Ricci curved manifolds. Yau’s work has caused a classical problem: does Picard’s theorem hold on complete Kähler manifolds with non-negative Ricci curvature? It is well-known that Nevanlinna theory [30] is a great development of both Liouville’s theorem and Picard’s theorem. So, a further problem is, can Nevanlinna theory extend to complete Kähler manifolds with non-negative Ricci curvature? As far as we know, neither of them has been solved.
Let be a complete non-compact Kähler manifold with non-negative Ricci curvature (see such examples in Sha-Yang [38] and Tian-Yau [43, 44]). The classical Picard’s problem states that
Picard’s Problem. Is every meromorphic function on necessarily a constant if it omits distinct values
Remark. Let be the universal covering, where is the unit disc in Thanks to the Liouville’s theorem of Yau, we may have an affirmative answer to the Picard’s problem if every holomorphic mapping can lift to a holomorphic mapping by By the mapping lifting theorem, admits a lifting if and only if
for some such that Since is simply-connected, then the fundamental group is trivial. It implies that is necessarily a trivial group if such a lifting exists. However, is not trivial in most cases. Thus, we cannot lift to in general.
An early result in this direction could be dated back to 1970, S. Kobayashi [24] obtained some Picard-type theorems for holomorphic mappings between complex manifolds. In particular, he showed that
Theorem A (Kobayashi, 1970).
If a complex Lie group acts on transitively, then every meromorphic function on must be a constant if it omits distinct values.
Kobayashi’s result is subject to certain complex manifolds which are acted on transitively by a complex Lie group. In 1975, Goldberg-Ishihara-Petridis [21] established some Picard-type theorems on the locally flat manifolds (see N. Petridis [31] also). They showed that
Theorem B (Goldberg-Ishihara-Petridis, 1975).
If is locally flat, then every holomorphic mapping of bounded dilatation must be a constant mapping.
In 2010, A. Atsuji [2] gave an affirmative answer for a class of meromorphic functions of slow growth rate by developing a Nevanlinna-type theory based on heat diffusion, Let us introduce his main result. Let be the Kähler form and Laplace-Beltrami operator of associated to respectively. One can endow with the Fubini-Study metric Let be a meromorphic function. The Kählerness of implies the Hilbert-Schmidt norm of the differential (with respect to ) can be formulated as
He showed that
Theorem C (Atsuji, 2010).
Let be a meromorphic function. Assume that satisfies the growth condition
for any where is the geodesic ball centered at a fixed reference point with radius in Then, must be a constant if it omits distinct values.
Atsuji’s method [2] (see [12] also) is using stochastic calculus by Brownian motion, who introduced the so-called Nevanlinna-type functions and where is the time of the Brownian motion on By computing curvatures and using Itô’s formula (see [22]), he established a heat diffusion version of the Second Main Theorem which leads to the above Theorem C. Note that is stochastically complete (i.e., is conservative), since has non-negative Ricci curvature due to Grigor’yan’s criterion [17]. However, in order to make and meaningful in the sense of Nevanlinna theory, the following conditions are necessary:
for ;
as ;
if omits
To do so, has to satisfy certain growth assumptions. That is why a growth condition is needed in Atsuji’s theorem. Unfortunately, the growth condition seems so strong that there are only a few meromorphic functions satisfy it.
The main purpose of the present paper is to generalize Nevanlinna theory to complete Kähler manifolds with non-negative Ricci curvature. It is known that Nevanlinna theory in several complex variables (see [28, 33]) studies the value distribution of meromorphic mappings between complex spaces, which has been developed for a long time. During this period, fruitful results have been achieved. For example, the reader may refer to L. Ahlfors [5], H. Cartan [7], Carlson-Griffths-King [8, 19], J. Noguchi [27, 28], E. Nochka [29], M. Ru [32, 33], B. Shiffman [34], B. Shabat [35], F. Sakai [36, 37], W. Stoll [41, 42], P. Vojta [40] and H. Wu [47], etc.; and also refer to [1, 2, 3, 12, 13, 14, 15]. To our knowledge, almost all results in Nevanlinna theory are built on complex manifolds with a complete Kähler metric of non-positive sectional curvature, while we seem to know little if the domain manifolds are not non-positively curved, since it is difficult to estimate the local Green function for a general bounded domain, even for a geodesic ball, under other curvature conditions.
Recently, the author [15] gave the first result on the Second Main Theorem in Nevanlinna theory on complete Kähler manifolds with non-negative Ricci curvature by introducing the global Green function method. Consequently, the author answered the Picard’s problem affirmatively provided that is of maximal volume growth:
where denotes the Riemannian volume of a geodesic ball centered at a fixed reference point with radius in
The original technique in [15] is to construct a family of relatively compact exhaustive domains of using the minimal positive global Green function for Based on an asymptotic estimate of obtained by Colding-Minicozzi [9], one can obtain two-sided bounds of the local Green function for in terms of integral forms. As a result, the author established the Second Main Theorem of meromorphic mappings on Let us introduce his main result as follows.
Colding-Minicozzi [9] (see [25] also) described the asymptotic behavior of : there exists a constant such that
where denotes the Riemannian distance function of from By means of the asymptotic estimate, the author defined a family of relatively compact domains as follows
for Hence, the Green function of for with a pole at satisfying Dirichlet boundary condition can be formulated as
According to Lemma 3.5 in [15], for any there exists such that
holds for all satisfying Based on this estimate, the Calculus Lemma was established. For is conformally equivalent to since has non-negative Ricci curvature and maximal volume growth (see [25]). So, by employing the standard arguments, the author established the Second Main Theorem of meromorphic mappings from into a complex projective manifold (see [15], Theorem 3.10), which gives an immediate result:
Theorem D (Dong, 2023).
If is of maximal volume growth, then every meromorphic function on must be a constant if it omits distinct values.
However, the technique in [15] depends on Colding-Minicozzi’s asymptotic estimate. Unfortunately, there are no such asymptotic estimates for if is not of maximal volume growth.
In this paper, we will generalize the Nevanlinna theory in the case that is a non-parabolic manifold, i.e., there exists a positive global Green function for Note that is non-parabolic if is of maximal volume growth. By N. Varopoulos [45, 46] (see Li-Tam-Wang [23, 25] also), the non-parabolicity implies that there exists a unique minimal positive global Green function of for denoted also by Our original idea is the construction of domain and application of Li-Yau’s estimate on heat kernel and global Green function. Using Li-Yau’s estimate [26], there exist constants such that
holds for all and For we redefine
Then, we have
which defines the Green function of for with a pole at satisfying Dirichlet boundary condition. There is a natural relationship between and : for we have
which plays an important role in the establishment of Nevanlinna theory in this paper (see Lemma 4.1).
Negative Curvature Method and Logarithmic Derivative Lemma Method are very typical approaches in the study of Nevanlinna theory. In this paper, the Logarithmic Derivative Lemma Method is used. So, we need to establish the so-called Calculus Lemma which is a way to the Logarithmic Derivative Lemma. We will introduce two techniques in the paper: Gradient Estimates (see Theorem 4.2) and Calculus Lemma (see Theorem 5.2) as follows.
Gradient Estimates. For any there exists such that
holds for all
The above gradient estimate yields an estimate of the harmonic measure on with respect to : for any there exists such that
holds for all (see Corollary 4.3).
Calculus Lemma. Let be a locally integrable function on Assume that is locally bounded at Then for any there exist a constant and a subset of finite Lebesgue measure such that
holds for all outside
Let be a nonconstant meromorphic function on The Calculus Lemma can derive the Logarithmic Derivative Lemma: for any there exists a subset of finite Lebesgue measure such that
holds for all outside (see Theorem 5.6).
1.2. Main results
Let be a complex projective manifold with over which one can put a Hermitian positive line bundle with Chern form We now take a divisor where denotes the complete linear system of Given a meromorphic mapping we have the Nevanlinna’s functions and (see Section 3.2). Let be the Ricci form of associated with the metric and be the canonical line bundle over The characteristic function of is defined by
The main result is the Second Main Theorem as follows.
Theorem 1.1 (=Theorem 6.1).
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature. Let be a complex projective manifold of complex dimension not greater than that of Let be a reduced divisor of simple normal crossing type, where is a positive line bundle over Let be a differentiably non-degenerate meromorphic mapping. Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
For a divisor the simple defect of with respect to is defined by
Set
Since then we obtain a defect relation:
Corollary 1.2 (=Corollary 6.3).
Assume the same conditions as in Theorem Then
Let denote the scalar curvature of Set
In the non-parabolic case, we answer the Picard’s problem as follows.
Corollary 1.3 (=Corollary 6.10).
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature. Every meromorphic function on reduces to a constant if it omits distinct values, where denotes the maximal integer not greater than
It implies that every meromorphic function with on reduces to a constant if it omits 2 distinct values.
2. Some Facts from Geometric Analysis
2.1. Volume Comparison Theorem
A space form is a complete Riemannian manifold with constant sectional curvature. Let be an -dimensional (simply-connected) space form with constant sectional curvature Denote by the Riemannian volume of a geodesic ball with radius in Let be an -dimensional complete Riemannian manifold with Ricci curvature denoted by Denote by the Riemannian volume of the geodesic ball centered at a fixed point with radius in
The well-known volume comparison theorem by Bishop-Gromov (see [6]) states that
Theorem 2.1 (Volume Comparison Theorem).
If for a constant then the volume ratio is non-increasing in and it tends to as Hence, we have
holds for all
In particular, if then we obtain:
Corollary 2.2.
We have
holds for all where is the volume of the unit ball in
When Calabi-Yau (see [39]) gave a lower bound of
Theorem 2.3.
Assume that is non-compact. If then has an infinite volume. More precisely, for any there exists a constant such that
holds for all
2.2. Estimates of Heat Kernels
Let be a complete Riemannian manifold of dimension with Laplace-Beltrami operator Fix a reference point We denote by the Riemannian distance function of from The heat kernel of is the minimal positive fundamental solution to the following heat equation
Theorem 2.4.
Assume that has non-negative Ricci curvature. Then for any there exist constants and such that
holds for all and
Set
This infinite integral is convergent if only if is non-parabolic. When is non-parabolic, is the unique minimal positive global Green function of for with a pole at i.e.,
where is the Dirac function with a pole at
When is non-parabolic with non-negative Ricci curvature, Li-Yau [26] (see [39] also) obtained an estimate of as follows.
Theorem 2.5.
Assume that is non-parabolic. If has non-negative Ricci curvature, then there exist constants depending only on such that
holds for all
3. Nevanlinna’s Functions and First Main Theorem
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature, of complex dimension The Kähler form of associated to the Kähler metric is defined by
3.1. Construction of
Fix a reference point Let be the Riemannian volume of where is the geodesic ball centered at with radius in Note that the non-parabolicity of implies that
Thus, we have the unique minimal positive global Green function of the half Laplace-Beltrami operator for which can be written as
where is the heat kernel of Let be the Riemannian distance function of from Using Theorem 2.5, there exist constants such that
holds for all For define
By
we see that is a bounded domain containing such that
Furthermore, the family exhausts i.e., for any sequence with we have
Whence, the boundary of can be formulated as
Due to Sard’s theorem, is a submanifold of for almost all
Set
Evidently, is the positive Green function of for with a pole at satisfying Dirichelet boundary condition, i.e.,
Let denote the harmonic measure on with respect to defined by
where is the inward normal derivative on is the Riemannian area element of
3.2. Nevanlinna’s Functions
In what follows, we will introduce Nevanlinna’s functions. Let be a meromorphic mapping, where is a complex projective manifold. Let be a Hermitian holomorphic line bundle over with the Chern form where
Fix a divisor where is the complete linear system of Let be the canonical section associated to i.e., is a holomorphic section of over with zero divisor The characteristic function, proximity function, counting function and simple counting function of are respectively defined by
where is the Riemannian volume element of
Equip with the Kähler metric Let be the Hilbert-Schmidt norm of differential with respect to the metrics We have
Locally, write where is a holomorphic local frame of and is a holomorphic function. By Poincaré-Lelong formula (see, e.g., [8]), it leads to
in the sense of currents. So, we obtain the alternative expressions of and as follows
and
3.3. First Main Theorem
Lemma 3.1 (Jensen-Dynkin formula).
Let be a -class function on outside a polar set of singularities at most. Assume that Then
Assume that Apply Jensen-Dynkin formula to then we are led to
Therefore, we are led to that
Theorem 3.2 (First Main Theorem).
Assume that Then
4. Gradient Estimates of Green Functions
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature, of complex dimension
Lemma 4.1.
We have
holds for all where is given by
Proof.
According to the definition of Green function for it is immediate that for
Since
then we obtain
Let denote the gradient operator on Then, we have an upper bound of as follows.
Theorem 4.2.
For any there exists such that
holds for all where is given by
Proof.
It yields from (1) that
Since
then the L’Hospital’rule yields that
where is the cut locus of and is the inward normal derivative on By
we see that for any there exists such that
holds for all Since has measure 0, with the aid of the continuity of on and on one then has the theorem proved. ∎
As a result, we obtain an estimate for upper bounds of harmonic measure on with respect to as follows:
Corollary 4.3.
For any there exists such that
holds for all where is given by
5. Calculus Lemma and Logarithmic Derivative Lemma
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature, of complex dimension
5.1. Calculus Lemma
Lemma 5.1 (Borel’s Lemma).
Let be a non-decreasing function on with Then for any there exists a subset of finite Lebesgue measure such that such that
holds for all outside
Proof.
The conclusion is clearly true for Next, we assume that Since is a non-decreasing function, then there exists a number such that The non-decreasing property of implies that the limit
exists, here is allowed. If then Set
Note that exists for almost every Then, we have
This completes the proof. ∎
We establish the Calculus Lemma as follows.
Theorem 5.2 (Calculus Lemma).
Let be a locally integrable function on Assume that is locally bounded at Then for any there exist a constant and a subset of finite Lebesgue measure such that
holds for all outside
Proof.
Invoking Lemma 4.1, we have
Set
A simple computation leads to
In further, we have
Apply Borel’s lemma to the left hand side of the above equality twice: one is to and the other is to we see that for any there exists a subset of finite Lebesgue measure such that
holds for all outside On the other hand, Corollary 4.3 implies that there exists such that
holds for all where is given by (1). Combining the above, we have
holds for all outside Therefore, we have the theorem proved by setting ∎
Corollary 5.3.
Let be a locally integrable function on Assume that is locally bounded at Assume that is of maximal volume growth. Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
5.2. Logarithmic Derivative Lemma
Let be a meromorphic function on The norm of the gradient is defined by
in a holomorphic local coordinate where is the inverse of Again, define the Nevanlinna’s characteristic function of by
where
It is not difficult to show that
On one puts a singular metric
which leads to
Lemma 5.4.
We have
Proof.
Note that
Whence, it concludes from Fubini’s theorem that
Lemma 5.5.
Let be a meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
Proof.
Define
Next, we establish the following Logarithmic Derivative Lemma.
Theorem 5.6 (Logarithmic Derivative Lemma).
Let be a meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
Proof.
Note that
By Lemma 5.5, we have for any there exists a subset of finite Lebesgue measure such that
holds for all outside ∎
By we have:
Corollary 5.7.
Let be a meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
6. Second Main Theorem and Defect Relation
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature, of complex dimension Let be complex projective manifold of complex dimension not greater than Over one can put a positive holomorphic line bundle Let be a reduced divisor of simple normal crossing type. Recall that
where is the Ricci form of
Let be a divisor, where are prime divisors. The reduced form of is defined by
6.1. Second Main Theorem
Write as the irreducible decomposition of Equiping every holomorphic line bundle with a Hermitian metric such that it induces the Hermitian metric on Pick such that with On we define a singular volume form
where
Let be a differentiably non-degenerate meromorphic mapping. Set
It is clear that
In further, we have
in the sense of currents. Thus, it yields that
We establish a Second Main Theorem as follows.
Theorem 6.1 (Second Main Theorem).
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature. Let be a complex projective manifold of complex dimension not greater than that of Let be a reduced divisor of simple normal crossing type, where is a positive line bundle over Let be a differentiably non-degenerate meromorphic mapping. Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
Proof.
Since is of simple normal crossing type, then we can provide a finite open covering of and finitely many rational functions on for every such that are holomorphic on with that
Moreover, we may require for On write
where is a positive smooth function on Let be a partition of the unity subordinate to Set
Again, put On we have
Fix any we can choose a holomorphic local coordinate near and a holomorphic local coordinate near such that
Set
Then, we have and at :
Define a non-negative function on by
Moreover, put for Then, at :
Whence, at :
Put together the above, we are led to
on Since is bounded on and
then we obtain
By Lemma 3.1
Combining (5) with (6) and using Theorem 5.6, we have
Using Theorem 5.2 and (4), for any there exists a subset of finite Lebesgue measure such that
holds for all outside Hence, we conclude that
for all outside By this with (6.1), we prove the theorem. ∎
Corollary 6.2.
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature. Let be a complex projective manifold of complex dimension not greater than that of Let be a reduced divisor of simple normal crossing type, where is a positive line bundle over Let be a differentiably non-degenerate meromorphic mapping. Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
6.2. Defect Relation
The simple defect of with respect to is defined by
By the First Main Theorem, we have
By the use of Theorem it is immediate that
Corollary 6.3 (Defect Relation).
Assume the same conditions as in Theorem Then
Let
denote the Ricci curvature tensor of Since is Kählerian, then
Again, denote by the scalar curvature of Note that is the trace of thus we have
It yields that
Corollary 6.4.
Let be a meromorphic mapping. If there exists a number such that satisfies the growth condition
for sufficiently large then must be differentiably degenerate.
Proof.
Assume on the contrary that is differentiably non-degenerate. Then, we obtain as The growth condition implies that
Thus, we deduce due to Corollary 6.3, but which contradicts with ∎
Consider a compact Riemann surface of genus Note that
for an arbitrary point where is the holomorphic line bundle over defined by Combining this with Theorem 6.1, we conclude that
Corollary 6.5.
Let be a nonconstant meromorphic mapping. Let be distinct points in Then
Let
be a meromorphic mapping. Then
Let be the hyperplane line bundle over Since
then Theorem 6.1 yields that
Corollary 6.6.
Let be a differentiably non-degenerate meromorphic mapping. Let be hyperplanes in general position in Then
A meromorphic function on can be seen as a meromorphic mapping Set
We obtain a Picard’s theorem as follows.
Corollary 6.7 (Picard’s Theorem).
Every meromorphic function on must be a constant if it omits distinct values, where denotes the maximal integer not greater than
In the following, we shall derive a better Picard’s theorem compared with Corollary 6.7. On we introduce the spherical distance
Let be a meromorphic function on which is regarded as a meromorphic mapping
In terms of the spherical distance, the proximity function of with respect to can be defined as
Adopt the Ahlfors-Shimizu’s characteristic function of as follows
Define
Then, we have the First Main Theorem
Theorem 6.8 (Second Main Theorem).
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature. Let be distinct values in Let be a nonconstant meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
Proof.
Define
Set
Clearly, we see that
Corollary 6.9 (Defect Relation).
Assume the same conditions as in Theorem Then
We derive a better Picard’s theorem as follows.
Corollary 6.10 (Picard’s Theorem).
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature. Every meromorphic function on must be a constant if it omits distinct values, where denotes the maximal integer not greater than