### v7

# On Picard’s Problem via Nevanlinna Theory

Abstract.
We consider the classical Picard’s problem for non-parabolic complete Kähler manifolds with non-negative Ricci curvature. Based on the global Green function approach, we give a positive answer to Picard’s problem under certain condition by developing Nevanlinna theory. That is, we prove that every meromorphic function on such a manifold reduces to a constant if it omits three distinct values, provided that the manifold satisfies a volume growth condition.
Key words and phrases:
Nevanlinna theory; value distribution; Picard’s theorem; defect relation; Kähler manifolds2010 Mathematics Subject Classification:
32H30; 32H25; 32A221. Introduction
1.1. Motivation
The famous Picard’s theorem asserts that a meromorphic function on the complex Euclidean spaces must reduce to a constant if it omits three distinct values. Historically, many authors have been committed to extensions of this theorem for a long time (see [4, 13, 15, 17, 18, 21, 28]). The study of Picard’s problem on a manifold may have originated from the celebrated work due to S. T. Yau [46] on Liouville’s theorem, which proves the Liouville’s property of harmonic functions on Ricci non-negatively curved manifolds. Yau’s work inspired people to think about the following Picard’s problem. Let be a complete non-compact Kähler manifold with non-negative Ricci curvature (see examples in Sha-Yang [35] and Tian-Yau [41, 42]). The classical Picard’s problem asks that
Picard’s Problem. Is every meromorphic function on necessarily a constant if it omits distinct values
Remark 1. Let be the universal covering, where is the unit disc in Thanks to the Liouville’s theorem by Yau, one may give an affirmative answer to the Picard’s problem if every holomorphic mapping can lift to a holomorphic mapping by Using the mapping lifting theorem, admits a lifting if and only if
for some such that Since is simply-connected, then the fundamental group is trivial. It implies that is necessarily a trivial group if such a lifting exists. However, is not trivial in most cases. So, we cannot lift to in general.
An early result in this direction could be dated back to 1970, S. Kobayashi [21] obtained some Picard-type theorems for holomorphic mappings between certain complex manifolds. In particular, he showed that
Theorem A (Kobayashi, 1970).
If a complex Lie group acts on transitively, then every meromorphic function on must be a constant if it omits distinct values.
Kobayashi’s result is subject to certain complex manifolds which are acted on transitively by a complex Lie group. In 1975, Goldberg-Ishihara-Petridis [18] treated a class of locally flat manifolds (see N. Petridis [28] also). They showed that
Theorem B (Goldberg-Ishihara-Petridis, 1975).
If is locally flat, then every holomorphic mapping of bounded dilatation must be a constant.
It is known that Nevanlinna theory [27] is a great development of Picard’s theorem, which investigates the value distribution of meromorphic mappings between complex manifolds. Refer to L. Ahlfors [5], H. Cartan [7], Carlson-Griffths-King [8, 16], J. Noguchi [24, 25], E. Nochka [26], M. Ru [29, 30], B. Shiffman [31], B. Shabat [32], F. Sakai [33, 34], W. Stoll [39, 40], P. Vojta [38] and H. Wu [45], and see [1, 2, 3, 10, 11, 12] and references therein.
In 2010, A. Atsuji [2] gave an affirmative answer for a class of meromorphic functions of slow growth rate by establishing a Nevanlinna-type theory based on heat diffusion, We shall introduce his main work. Let be the Kähler form and Laplace-Beltrami operator of associated to respectively. One can endow with the Fubini-Study metric Let be a meromorphic function. The Kählerness of implies the Hilbert-Schmidt norm of the differential (with respect to ) can be formulated as
He showed that
Theorem C (Atsuji, 2010).
Let be a meromorphic function. Assume that satisfies the growth condition
for any where is the geodesic ball centered at a fixed reference point with radius in Then, must be a constant if it omits distinct values.
Atsuji’s method [2] (see [10] also) is to use stochastic calculus by Brownian motion. Atsuji successfully introduced Nevanlinna-type functions and where is the time of the Brownian motion on By computing curvatures and utilizing Itô’s formula (see [19]), he established an analogue of Nevanlinna theory. Note that is stochastically complete (i.e., is conservative) since Grigor’yan’s criterion [14]. To make and meaningful, the following conditions are necessary:
for ;
as ;
if omits
To do so, has to satisfy certain growth assumption. That is why a growth condition is needed in Theorem C.
In this paper, we generalize Nevanlinna theory to non-parabolic complete Kähler manifolds with non-negative Ricci curvature. Our innovative method is to construct a family of exhaustive precompact domains for by means of a unique minimal positive global Green function for so that we can well define the Nevanlinna’s functions on and establish a Second Main Theorem by introducing the global Green function method. Apply the Second Main Theorem, we can give a positive answer to the Picard’s problem under certain condition.
1.2. Main results
Assume that is non-parabolic, by which we mean that there exists a positive global Green function for By N. Varopoulos [43, 44] (see Li-Tam-Wang [20, 22] also), the non-parabolicity implies that there exists a unique minimal positive global Green function of for denoted by where is a fixed reference point in Using Li-Yau’s estimate [23], there are constants such that
holds for all and where is the Riemannian distance function of from and denotes the Riemannian volume of the geodesic ball centered at with radius in For define
Then, for all and the family exhausts
Let be a complex projective manifold with over which one can put a Hermitian positive line bundle with Chern form We now take a divisor where denotes the complete linear system of Given a meromorphic mapping we have the Nevanlinna’s functions and on (see Section 3.2). Let denote the Ricci form of associated to the metric and be the canonical line bundle over The characteristic function of is defined by
Again, set
The main result is the following Second Main Theorem.
Theorem 1.1 (=Theorem 6.1).
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature. Let be a complex projective manifold of complex dimension not greater than that of Let be a reduced divisor of simple normal crossing type, where is a positive line bundle over Let be a differentiably non-degenerate meromorphic mapping. Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside where is given by
For a divisor the simple defect of with respect to is defined by
Set
We obtain a defect relation:
Corollary 1.2 (=Corollary 6.2).
Assume the same conditions as in Theorem Then
if one of the following conditions is satisfied
satisfies the volume growth condition
is of non-polynomial type growth, i.e., satisfies the growth condition
Remark 2. Note that due to
The condition (2) is relaxed. For example, it holds when with and holds when with or and
As an application of Corollary 1.2, we answer the Picard’s problem under certain condition as follows.
Corollary 1.3.
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature. If satisfies the volume growth condition then every meromorphic function on reduces to a constant if it omits distinct values.
Corollary 1.4.
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature. Then, every meromorphic function of non-polynomial type growth on can omit distinct values at most.
2. Some Facts from Geometric Analysis
2.1. Volume Comparison Theorem
A space form is a complete Riemannian manifold with constant sectional curvature. Let be an -dimensional (simply-connected) space form with constant sectional curvature Denote by the Riemannian volume of a geodesic ball with radius in Let be an -dimensional complete Riemannian manifold with Ricci curvature denoted by Denote by the Riemannian volume of the geodesic ball centered at a fixed point with radius in
The well-known volume comparison theorem by Bishop-Gromov (see [6]) states that
Theorem 2.1.
If for a constant then the volume ratio is non-increasing in and it tends to as Hence, we have
holds for all
When we obtain:
Corollary 2.2.
We have
holds for all where is the volume of the unit ball in
When Calabi-Yau (see [36]) proved that
Theorem 2.3.
Assume that is non-compact. If then has an infinite volume. More precisely, for any there exists a constant such that
holds for all
2.2. Estimates of Heat Kernel
Let be a complete Riemannian manifold of dimension with Laplace-Beltrami operator Fix a reference point We denote by the Riemannian distance function of from The heat kernel of is the minimal positive fundamental solution to the following heat equation
Theorem 2.4.
Assume that has non-negative Ricci curvature. Then for any there exist constants and such that
holds for all and
Set
This infinite integral is convergent if only if is non-parabolic. When is non-parabolic, is the unique minimal positive global Green function of for with a pole at i.e.,
where is the Dirac’s delta function with a pole at
When is non-parabolic with non-negative Ricci curvature, Li-Yau [23] (see [36] also) gave two-sided bounds of as follows.
Theorem 2.5.
Assume that is non-parabolic. If has non-negative Ricci curvature, then there exist constants depending only on such that
holds for all
3. Nevanlinna’s Functions and First Main Theorem
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature, of complex dimension The Kähler form of associated to the Kähler metric is defined by
3.1. Construction of
Fix a reference point Let be the Riemannian volume of where is the geodesic ball centered at with radius in Note that the non-parabolicity of implies that
Thus, we have the unique minimal positive global Green function of the half Laplace-Beltrami operator for which can be written as
where is the heat kernel of Let be the Riemannian distance function of from Using Theorem 2.5, there exist constants such that
holds for all For define
By
where means we derive that is a precompact domain containing such that
In fact, the family exhausts namely, for any sequence with we have
Whence, the boundary of can be formulated as
Due to Sard’s theorem, is a submanifold of for almost all
Set
Evidently, is the positive Green function of for with a pole at satisfying Dirichelet boundary condition, i.e.,
Let denote the harmonic measure on with respect to defined by
where is the inward normal derivative on is the Riemannian area element of
3.2. Nevanlinna’s Functions
In what follows, we will introduce Nevanlinna’s functions. Let be a meromorphic mapping, where is a complex projective manifold. Let be a Hermitian holomorphic line bundle over with the Chern form where
Fix a divisor where is the complete linear system of Let be the canonical section associated to i.e., is a holomorphic section of over with zero divisor The characteristic function, proximity function, counting function and simple counting function of are respectively defined by
where is the Riemannian volume element of
Locally, write where is a holomorphic local frame of and is a holomorphic function. By Poincaré-Lelong formula (see, e.g., [8]), it leads to
in the sense of currents. So, we obtain the alternative expressions of and as follows
and
3.3. First Main Theorem
Lemma 3.1 (Jensen-Dynkin formula).
Let be a -class function on outside a polar set of singularities at most. Assume that Then
Assume that Apply Jensen-Dynkin formula to then we are led to
Therefore, we are led to that
Theorem 3.2 (First Main Theorem).
Assume that Then
4. Gradient Estimates of Green Functions
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature, of complex dimension
Lemma 4.1.
We have
holds for all where is given by
Proof.
According to the definition of Green function for it is immediate that for
Since
we obtain
Let denote the gradient operator on a Riemannian manifold. Cheng-Yau [9] proved the following theorem.
Lemma 4.2.
Let be a complete Riemannian manifold. Let be a non-negative harmonic function on a geodesic ball centered at with radius Then, there exists a constant depending only on the dimension of such that
where is the Riemannian distance between and and is the lower bound of the Ricci curvature on
We have an upper bound of as follows.
Theorem 4.3.
There exists a constant such that
Proof.
By
we can obtain an estimate for upper bounds of as follows.
Corollary 4.4.
There exists a constant such that
where is the Riemannian area element of
5. Calculus Lemma and Logarithmic Derivative Lemma
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature, of complex dimension
5.1. Calculus Lemma
Lemma 5.1 (Borel’s Lemma).
Let be a non-decreasing function on with Then for any there exists a subset of finite Lebesgue measure such that such that
holds for all outside
Proof.
The conclusion is clearly true for Next, we assume that Since is a non-decreasing function, then there exists a number such that The non-decreasing property of implies that the limit
exists, here is allowed. If then Set
Note that exists for almost all Then, we have
This completes the proof. ∎
We establish the Calculus Lemma as follows.
Theorem 5.2 (Calculus Lemma).
Let be a locally integrable function on Assume that is locally bounded at Then there exists a constant such that for any there exists a subset of finite Lebesgue measure such that
holds for all outside where is given by
Proof.
Invoking Lemma 4.1, we have
Set
A simple computation leads to
In further, we have
Apply Borel’s lemma to the left hand side of the above equality twice: one is to and the other is to then we deduce that for any there exists a subset of finite Lebesgue measure such that
holds for all outside On the other hand, Corollary 4.4 implies that there exists a constant such that
Set Combining the above, we have
holds for all outside ∎
5.2. Logarithmic Derivative Lemma
Let be a meromorphic function on The norm of the gradient is defined by
in holomorphic local coordinates here is the inverse of Again, define the Nevanlinna’s characteristic function of by
where
It is not difficult to show that
On one puts a singular metric
which leads to
Lemma 5.3.
We have
Proof.
Note that
Whence, it concludes from Fubini’s theorem that
Lemma 5.4.
Let be a meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside where is given by
Proof.
Define
We establish the following Logarithmic Derivative Lemma.
Theorem 5.5 (Logarithmic Derivative Lemma).
Let be a meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside where is given by
Proof.
Note that
By Lemma 5.4, we have for any there exists a subset of finite Lebesgue measure such that
holds for all outside where is given by ∎
6. Second Main Theorem and Defect Relation
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature, of complex dimension Let be complex projective manifold of complex dimension not greater than Over one can put a positive holomorphic line bundle Let be a reduced divisor of simple normal crossing type. Recall that
where is the Ricci form of
Let be a divisor, where are prime divisors. The reduced form of is defined by
Write as the irreducible decomposition of Equiping every holomorphic line bundle with a Hermitian metric such that it induces the Hermitian metric on Pick such that with On we define a singular volume form
where
Let be a differentiably non-degenerate meromorphic mapping. Set
It is clear that
In further, we have
in the sense of currents. Thus, it yields that
We establish a Second Main Theorem as follows.
Theorem 6.1 (Second Main Theorem).
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature. Let be a complex projective manifold of complex dimension not greater than that of Let be a reduced divisor of simple normal crossing type, where is a positive line bundle over Let be a differentiably non-degenerate meromorphic mapping. Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside where is given by
Proof.
Since is of simple normal crossing type, then there are a finite open covering of and finitely many rational functions on for each such that are holomorphic on with that
In addition, we can require that for all On write
where is a positive smooth function on Let be a partition of the unity subordinate to Set
Again, put On we have
Fix any one can take holomorphic local coordinates near and holomorphic local coordinates near such that
Set
Then, we have and at :
Define a non-negative function on by
Moreover, put for Then, at :
Whence, at :
Put together the above, we are led to
on Since is bounded on and
then we obtain
By Lemma 3.1
Combining (6) with (7) and using Theorem 5.5, we have
Using Theorem 5.2 and (5), for any there exists a subset of finite Lebesgue measure such that
holds for all outside Hence, we conclude that
for all outside By this with (6), we prove the theorem. ∎
Recall that the simple defect of with respect to is defined by
By the First Main Theorem, we have
By the use of Theorem it is immediate that
Corollary 6.2 (Defect Relation).
Assume the same conditions as in Theorem Then
if one of the following conditions is satisfied
satisfies the volume growth condition
is of non-polynomial type growth, i.e., satisfies the growth condition
Proof.
Note that due to Corollary 2.2. We have
Using the volume growth condition, we obtain
Since a non-constant meromorphic mapping has the growth rate at least, then we have the theorem proved if is satisfied due to Theorem 6.1.
For we have
where
So, it yields from that
This completes the proof. ∎