### v8

# On Picard’s Problem via Nevanlinna Theory

Abstract.
We consider the classical Picard’s problem for non-parabolic complete Kähler manifolds with non-negative Ricci curvature. Based on the global Green function approach, we give a positive answer to Picard’s problem under certain condition by developing Nevanlinna theory. That is, we prove that every meromorphic function on such a manifold reduces to a constant if it omits three distinct values, provided that the manifold satisfies a volume growth condition; and prove that every meromorphic function of non-polynomial type growth on such a manifold can avoid 2 distinct values at most.
Key words and phrases:
Nevanlinna theory; value distribution; Picard’s theorem; defect relation; Kähler manifolds2020 Mathematics Subject Classification:
32H30; 32H25; 32A221. Introduction
1.1. Motivation
The famous Picard’s theorem asserts that a meromorphic function on the complex Euclidean spaces must reduce to a constant if it omits three distinct values. Historically, many authors have been committed to extensions of this theorem for a long time (see [4, 13, 15, 17, 18, 21, 28]). The study of Picard’s problem on a manifold may have originated from the celebrated work due to S. T. Yau [47] on Liouville’s theorem, which proves the Liouville’s property of harmonic functions on Ricci non-negatively curved manifolds. Yau’s work inspires people to think about the following Picard’s problem. Let be a complete noncompact Kähler manifold with non-negative Ricci curvature (see examples for such manifolds in Sha-Yang [35] and Tian-Yau [40, 41]).
The classical Picard’s problem states that
Picard’s Problem. Is every meromorphic function on necessarily a constant if it omits distinct values
Remark 1. Let be the universal covering, where is the unit disc in Thanks to the Liouville’s theorem by Yau, one may give an affirmative answer to the Picard’s problem if every holomorphic mapping can lift to a holomorphic mapping by Using the mapping lifting theorem, admits a lifting if and only if
for some such that Since is simply-connected, the fundamental group is trivial. This implies that is necessarily a trivial group if such a lifting admits. However, is not trivial in most cases. Thus, we cannot lift to in general.
An early result in this direction could be dated back to 1970, S. Kobayashi [21] obtained some Picard-type theorems for holomorphic mappings between certain complex manifolds. In particular, he showed that
Theorem A (Kobayashi, 1970).
Assume that there is a complex Lie group acting on transitively. Then, every meromorphic function on must be a constant if it omits distinct values.
Kobayashi’s result is subject to certain complex manifolds which are acted on transitively by a complex Lie group. In 1975, Goldberg-Ishihara-Petridis [18] treated a class of locally flat manifolds (see N. Petridis [28] also). They showed that
Theorem B (Goldberg-Ishihara-Petridis, 1975).
If is locally flat, then every holomorphic mapping of bounded dilatation must be a constant.
It is known that Nevanlinna theory [27] is a great development of Picard’s theorem, which investigates the value distribution of meromorphic mappings between complex manifolds. For instance, refer to L. Ahlfors [5], H. Cartan [7], Carlson-Griffths-King [8, 16], J. Noguchi [24, 25], E. Nochka [26], M. Ru [29, 30], B. Shiffman [31], B. Shabat [32], F. Sakai [33, 34], W. Stoll [38, 39], P. Vojta [45], H. Wu [46], and also refer to [1, 2, 3, 10, 11, 12] and references therein.
In 2010, A. Atsuji [2] gave an affirmative answer for a class of meromorphic functions with slower growth by introducing a Nevanlinna-type theory based on heat diffusion, We introduce his main work. Let stand for the Kähler form and Laplace-Beltrami operator on respectively. Let be a meromorphic function. We equip with Fubini-Study metric The Kählerness of indicates that the Hilbert-Schmidt norm of differential with respect to metrics can be expressed as
He showed that
Theorem C (Atsuji, 2010).
Let be a meromorphic function. Assume that satisfies the growth condition
for any where is the geodesic ball centered at a fixed reference point with radius in Then, must be a constant if it omits distinct values.
Atsuji’s method [2] (see [10] also) is to use stochastic calculus by Brownian motion. Atsuji successfully introduced Nevanlinna-type functions and where is the time of the Brownian motion on By computing curvatures and using Itô’s formula (see [19]), he established an analogue of the Nevanlinna theory. Note that is stochastically complete (i.e., is conservative) due to Grigor’yan’s criterion (see [14]). In order to make and meaningful, the following are necessary:
for ;
as ;
if omits
Hence, needs to satisfy certain growth assumption. That is why a growth condition is added to Theorem C.
In this paper, our purpose is to study the Picard’s problem using the tool of Nevanlinna theory, i.e., we shall extend the classical Nevanlinna theory to non-parabolic complete Kähler manifolds with non-negative Ricci curvature, and apply the Second Main Theorem to establish a Picard’s theorem. Let us first recall the basic notions of the non-parabolicity and the Ricci curvature of Riemannian manifolds. Let be a complete Riemannian manifold. One says that is non-parabolic, if there exists a positive global Green function for and parabolic otherwise. Or equivalently, in a viewpoint of geometric analysis, is called non-parabolic if
for and parabolic if this infinite integral is divergent, where denotes the transition density function of the Brownian motion generated by the Laplace-Beltrami operator on and which is also called the heat kernel of Note that when is non-parabolic, defines the unique minimal positive global Green function of for More details may refer to [20, 22, 42, 43]. Let stand for the Riemannian curvature tensor of The Ricci curvature tensor of is defined by
for any with where is an orthonormal basis of The Ricci curvature of at in a direction is defined by
We say that is of non-negative Ricci curvature at if for each nonzero vector and that is of non-negative Ricci curvature, if it is of non-negative Ricci curvature at each Moreover, the sectional curvature of at along a section spanned by is defend by
From the definitions, the sign of the sectional curvature at a point determines the sign of the Ricci curvature at the same point.
To facilitate the construction of examples later for manifolds considered in this paper, we shall provide some criteria of non-parabolicity via the volume growth. A sharp necessary condition by N. Varopoulos [44] states that if is non-parabolic, then
where denotes the Riemannian volume of a geodesic ball centered at a fixed reference point with radius in However, (1) is far from sufficient, and a counterexample was constructed in [44]. The first major result for the sufficiency was due to N. Varopoulos [43] and Li-Yau [23]. Based on the heat kernel estimates, they proved that if has non-negative Ricci curvature and (1) is satisfied, then is non-parabolic. So, for with non-negative Ricci curvature, is non-parabolic if and only if (1) is satisfied.
Next, assume that is a non-parabolic complete noncompact Kähler manifold with non-negative Ricci curvature, of complex dimension Here, the non-parabolicity of is meant that is non-parabolic as a Riemannian manifold. The Chern-Ricci form of on is defined by
with Since is Kähler, is equivalent to Our innovative idea in the present paper is to construct a family of exhaustive precompact domains for by means of the minimal positive global Green function for so that it allows us to well define Nevanlinna’s functions on and then establish a Second Main Theorem by estimating the gradient of Green function for Applying the Second Main Theorem, we provide a positive answer to the Picard’s problem under certain condition.
1.2. Main results
Let be the minimal positive global Green function of for in which is the Laplace-Beltrami operator on and is a fixed reference point in Using Li-Yau’s estimate [23], there are constants such that
holds for all and all in which is the Riemannian distance between and and stands for the Riemannian volume of the geodesic ball centered at with radius in For define
Then, we have for all and the family exhausts
Let be a complex projective manifold with over which one can put a Hermitian positive line bundle with Chern form We now take a divisor where denotes the complete linear system of Given a meromorphic mapping we have the Nevanlinna’s functions on (see Section 3.2). Let be the canonical line bundle over The characteristic function of is defined by
where is the Kähler form of and is the Green function of for with a pole at satisfying Dirichlet boundary condition. Set
The main result is the following Second Main Theorem.
Theorem 1.1 (=Theorem 6.1).
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature. Let be a complex projective manifold of complex dimension not greater than that of Let be a reduced divisor of simple normal crossing type, where is a positive line bundle over Let be a differentiably non-degenerate meromorphic mapping. Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside where is given by
For a divisor the simple defect of with respect to is defined by
Set
We obtain a defect relation:
Corollary 1.2 (=Corollary 6.2).
Assume the same conditions as in Theorem Then
if one of the following conditions is satisfied
satisfies the volume growth condition
is of non-polynomial type growth, i.e., satisfies the growth condition
Remark 2. Note that as due to
The condition (3) is relaxed. For example, it holds when with and holds when with or and
As an application of Corollary 1.2, we consider the Picard’s problem under certain condition below. Treat the case when with the point line bundle, i.e., the hyperplane line bundle Take where are distinct points in If we put then A basic and well-known fact states that
which yields that
whenever or in Corollary 1.2 is satisfied. According to the arguments above, we conclude that
Corollary 1.3.
Let be a non-parabolic complete noncompact Kähler manifold with non-negative Ricci curvature. If satisfies the volume growth condition then every meromorphic function on reduces to a constant if it omits distinct values.
Corollary 1.4.
Let be a non-parabolic complete noncompact Kähler manifold with non-negative Ricci curvature. Then, every meromorphic function of non-polynomial type growth on can omit distinct values at most.
To close the section, we finally provide some typical examples for where Corollaries 1.3 and 1.4 apply. According to (1), we first notice that with is a non-parabolic complete noncompact Kähler manifold with zero Ricci curvature. Below, we give two nontrivial examples.
Example 1. For an integer we equip with Fubini-Study metric Note that is a complete Kähler manifold with positive sectional curvature (and thus with positive Ricci curvature). We consider the product manifold with which is equipped with the product metric induced from and the standard Euclidean metric on Because is a complete noncompact Kähler manifold, so is Since is flat, is clearly Ricci non-negatively curved under this product metric. Again, by the dimension condition that we deduce that is of the volume growth: as which leads to that is satisfied. Therefore, is a non-parabolic complete noncompact Kähler manifold with non-negative Ricci curvature. It is not difficult to see that there exist a lot of nonconstant meromorphic functions on For instance, we can take or where is any meromorphic function on and is any rational function on
Example 2. Treat a product manifold with and where denotes the -dimensional complex torus equipped with the metric inherited from the standard Euclidean metric on Note that is a complete noncompact Kähler manifold with zero Ricci curvature under the induced product metric. By the similar argument as in Example 1 also shows the non-parabolicity of The existence of nonconstant meromorphic functions on can be confirmed easily. For instance, for we can take or where is any meromorphic function on and is any elliptic function with a periodic lattice on
More general, can be taken as the product of and in which is any complete compact Kähler manifold with non-negative Ricci curvature, and is a complete noncompact Kähler manifold with non-negative Ricci curvature and volume growth satisfying (1). In particular, satisfies the desired conditions.
2. Some Facts from Geometric Analysis
2.1. Volume Comparison Theorem
A space form is a complete (simply-connected) Riemannian manifold with constant sectional curvature. Let denote the -dimensional space form with constant sectional curvature and let denote the Riemannian volume of a geodesic ball with radius in Let denote a complete Riemannian manifold of dimension with Ricci curvature tensor and let denote the Riemannian volume of a geodesic ball centered at a fixed reference point with radius in
The well-known volume comparison theorem by Bishop-Gromov (see [6]) states that
Theorem 2.1.
If then the volume ratio is a non-increasing function in and which tends to as Hence, we have
holds for all
When we obtain:
Corollary 2.2.
We have
holds for all where denotes the standard Euclidean volume of the unit ball in
When Calabi-Yau (see [36]) showed that
Theorem 2.3.
Assume that is noncompact. If then has an infinite volume. More precisely, for any there exists a constant such that
holds for all
2.2. Estimates of Heat Kernels
Let be a complete Riemannian manifold of dimension with Laplace-Beltrami operator Fix a reference point We denote by the Riemannian distance function of from The heat kernel of is the minimal positive fundamental solution to the following heat equation
Li-Yau [23] (see [36] also) obtained the two-sided estimates of
Theorem 2.4.
Assume that has non-negative Ricci curvature. Then for any there exist constants and such that
holds for all and all
Set
This infinite integral is convergent if only if is non-parabolic. When is non-parabolic, is the unique minimal positive global Green function of for with a pole at i.e.,
where is the Dirac’s delta function with a pole at
When is non-parabolic with non-negative Ricci curvature, Li-Yau [23] (see [36] also) gave two-sided bounds of as follows.
Theorem 2.5.
Assume that has non-negative Ricci curvature. If is non-parabolic, then there exist constants depending only on such that
holds for all
3. Nevanlinna’s Functions and First Main Theorem
Let be a non-parabolic complete noncompact Kähler manifold with non-negative Ricci curvature, of complex dimension The Kähler form of associated to metric is defined by
3.1. Construction of
Fix a reference point Let be the Riemannian volume of where denotes the geodesic ball centered at with radius in Note that the non-parabolicity of implies that
Thus, we have the unique minimal positive global Green function of the half Laplace-Beltrami operator for which can be written as
where is the heat kernel of Let be the Riemannian distance function of from Using Theorem 2.5, there exist constants such that
holds for all For define
Since
one can conclude immediately that is a precompact domain containing such that whenever and that
Thus, the family exhausts i.e., for any sequence with we have
The boundary of can be described as
By Sard’s theorem, is smooth for almost all
Set
Evidently, is the positive Green function of for with a pole at satisfying Dirichlet boundary condition, i.e.,
where is the Dirac’s delta function with a pole at Let stand for the harmonic measure on with respect to defined by
where is the inward normal derivative on and is the induced Riemannian area element of
3.2. Nevanlinna’s Functions
In what follows, we will introduce Nevanlinna’s functions. Let be a meromorphic mapping, where is a complex projective manifold. Let be a Hermitian holomorphic line bundle over with the Chern form
where
so that
Fix a divisor where is the complete linear system of Let be the canonical section associated to i.e., is a holomorphic section of over with zero divisor The characteristic function, proximity function, counting function and simple counting function of are respectively defined by
where is the Riemannian volume element of
Locally, write in which is a local holomorphic frame of and is a holomorphic function. Using Poincaré-Lelong formula (see, e.g., [8]), we get
in the sense of currents. So, we obtain the alternative expressions of and as follows
and
3.3. First Main Theorem
Lemma 3.1 (Green-Dynkin formula).
Let be a function on outside a polar set of singularities at most. Assume that Then
Assume that Apply Green-Dynkin formula to we are led to
To conclude, we obtain:
Theorem 3.2 (First Main Theorem).
Assume that Then
4. Gradient Estimates of Green Functions
Let be a non-parabolic complete noncompact Kähler manifold with non-negative Ricci curvature, of complex dimension
To give a gradient estimate of we need to introduce some lemmas as follows.
Lemma 4.1.
We have
holds for all where is given by
Proof.
According to the definition of Green function for it is immediate that for
Since for we obtain
Let denote the gradient operator on any Riemannian manifold. Cheng-Yau [9] proved the following theorem.
Lemma 4.2.
Let be a complete Riemannian manifold of dimension Let be any geodesic ball centered at with radius in Then, there exists a constant depending only on such that
holds for any non-negative harmonic function on where is the Riemannian distance between and and is the lower bound of Ricci curvature of
We obtain an upper estimate of as follows.
Theorem 4.3.
There exists a constant independent of such that
Proof.
By
we can derive an upper estimate of as follows.
Corollary 4.4.
There exists a constant independent of such that
where is the induced Riemannian area element of
5. Calculus Lemma and Logarithmic Derivative Lemma
Let be a non-parabolic complete noncompact Kähler manifold with non-negative Ricci curvature, of complex dimension
5.1. Calculus Lemma
Lemma 5.1 (Borel’s Growth Lemma).
Let be a non-decreasing function on with Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
Proof.
The conclusion is clearly true for Next, we assume that Since is non-decreasing, there is a number such that Using the non-decreasing property of the limit exists or If then Set
Note that exists for almost all Then, we have
This completes the proof. ∎
We establish the following Calculus Lemma.
Theorem 5.2 (Calculus Lemma).
Let be a locally integrable function on Assume that is locally bounded at Then there is a constant independent of such that for any there exists a subset of finite Lebesgue measure such that
holds for all outside where is given by
Proof.
Invoking Lemma 4.1, we have
Set
A simple computation leads to
In further, we have
Employing Borel’s growth lemma to the left hand side of the above equality twice: one is to and another is to then we conclude that for any there exists a subset of finite Lebesgue measure such that
holds for all outside On the other hand, Corollary 4.4 implies that there exists a constant independent of such that
Set Combining the above, we have
holds for all outside The proof is completed. ∎
5.2. Logarithmic Derivative Lemma
Let be a meromorphic function on The norm of the gradient is defined by
in a local holomorphic coordinate where is the inverse of Define the Nevanlinna’s characteristic function of by
where
For any it is not hard to deduce that
Put a singular metric on :
Lemma 5.3.
We have
Proof.
A direct computation gives
Whence, we conclude from Fubini’s theorem that
Lemma 5.4.
Let be a meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside where is given by
Proof.
Define
With the above preparations, we are to prove the following Logarithmic Derivative Lemma.
Theorem 5.5 (Logarithmic Derivative Lemma).
Let be a meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside where is given by
Proof.
The concavity of gives that
By Lemma 5.4, we have for any there exists a subset of finite Lebesgue measure such that
holds for all outside where is given by ∎
6. Second Main Theorem and Defect Relation
Let be a non-parabolic complete noncompact Kähler manifold with non-negative Ricci curvature, of complex dimension Let be a complex projective manifold of complex dimension not greater than Over one can put a positive holomorphic line bundle Let be a reduced divisor of simple normal crossing type. Recall that
where is the Chern-Ricci form of on
Let be a divisor, where are prime divisors. The reduced form of is defined by
Write as the irreducible decomposition of Equiping every holomorphic line bundle with a Hermitian metric such that it induces the Hermitian metric on Pick such that with On we define a singular volume form
Let be a differentiably non-degenerate meromorphic mapping. Set
It is clear that
In further, we have
in the sense of currents. Thus, it yields that
We give the main theorem in this paper, i.e., the Second Main Theorem as follows.
Theorem 6.1 (Second Main Theorem).
Let be a non-parabolic complete noncompact Kähler manifold with non-negative Ricci curvature. Let be a complex projective manifold of complex dimension not greater than that of Let be a reduced divisor of simple normal crossing type, where is a positive line bundle over Let be a differentiably non-degenerate meromorphic mapping. Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside where is given by
Proof.
Since has the simple normal crossing type, there exist a finite open covering of and finitely many rational functions on for each such that are holomorphic on with
In addition, we can require that for all On write
where is a positive smooth function on Let be a partition of the unity subordinate to Set
Again, put On we have
Fix any Take a local holomorphic coordinate near and a local holomorphic coordinate near such that
and
Set Then, we have and
Define a non-negative function on by
Moreover, put for Then
which yields that
Put together the above, we are led to
on Since is bounded on and
we obtain
By Lemma 3.1
Combining (7) with (8) and using Theorem 5.5, we have
Using Theorem 5.2 and (6), for any there exists a subset of finite Lebesgue measure such that
holds for all outside Hence, we conclude that
for all outside By this with (6), we prove the theorem. ∎
Recall that the simple defect of with respect to is defined by
By the First Main Theorem, we have
By the use of Theorem it is immediate that
Corollary 6.2 (Defect Relation).
Assume the same conditions as in Theorem Then
if one of the following conditions is satisfied
satisfies the volume growth condition
is of non-polynomial type growth, i.e., satisfies the growth condition
Proof.
We first show that the corollary holds if is satisfied. By Corollary 2.2, we have Thus, we are led to that
Using the condition we obtain
Since each non-constant meromorphic mapping has growth at least, we have the theorem proved if is satisfied due to Theorem 6.1. For we obtain
where
Thus, it yields from that
This completes the proof. ∎